Im writing a lab report and have done a ttest but dont know how to report the information. it is a two tailed t test with a critical value of p=0.05.I got the mean of 1.07 s.d 0.3 and 2.64 s.d 0.77 n=6 with a t test value from excel of 0.004. how would I write this 1.07±0.3 ml/min for sample 1 and 2.64 ±0.77 ml/min sample 2 n=6 with paired t-test = 0.004 ,p value p˂0.05 is this right
Answer #1

Since you are doing two tailed paired test that means the hypotheses will be

Ho: u1=u2

Ha: u1\nequ2

Where u1, u2 are the population mean for sample 1 and 2 respectively.

Since the t statistic calculated as

T test= 0.004 , then the associated p value will be much closer to 0.99, at degree of freedom of n-1= 6-1=5 hence

P value is > 0.05 this is right

Conclusion:

Since p value is greater than the level of significance that is 0.05 we fail to reject the null hypothesis and we do no have enough evidence to support the claim.

But one more thing I think you have written T test value instead of 0 value as 0.004 if it is so then ,

P value < 0.05 it is correct and hence we conclude that we have enough evidence to reject the null hypothesis and support the claim.

First check the data